# Flow Straighter and Faster: Efficient One-Step Generative Modeling via MeanFlow on Rectified Trajectories

### v1

@@ 5.2 Training Efficiency
We also compare estimated FLOPs. Although Re-Meanflow shows a smaller advantage under this metric, FLOPs alone do not fully reflect practical runtime. A substantial portion of our compute lies in the inference-only reflow stage, which can be executed on widely accessible inference-grade GPUs and runs significantly faster than training workloads with the same FLOP count. Training steps incur considerable overhead, such as data loading and batching, and rely on computationally expensive backward and JVP operations that typically require high-end training GPUs.
Takeaway. This highlights a key practical benefit of Re-Meanflow: by shifting most of the computational burden away from scarce high-end training GPUs and onto widely available inference-grade accelerators, Re-Meanflow substantially improves the accessibility and real-world feasibility of one-step flow distillation.
@@ 5.3 Ablation Study
Distance Truncation.
Table 2: Ablation on truncation strength on ImageNet-$512^{2}$. Percentage improvement is computed relative to the no-truncation baseline.
Truncation Strength FID $\downarrow$ Improvement (%)
No Truncation [8] 3.50
Top $5\%$ 3.10 $11.4\%$
[…]
Top $15\%$ 3.19 $8.9\%$
We evaluate different truncation thresholds for Re-Meanflow, ranging from $5\%$ to $15\%$, and report the resulting FID scores in Tab. 2. All truncation levels outperform the no-truncation baseline. The best result is achieved with a $10\%$ threshold, which reduces FID from $3.5$ to $3.03$.
As the threshold increases beyond $10\%$, performance begins to degrade (e.g., $3.19$ at $15\%$). This pattern reflects the trade-off inherent in truncation: while discarding extremely long-distance pairs helps suppress the high-curvature tail, removing too many couplings reduces data coverage and weakens the supervision signal. Overall, the empirical trend supports our analysis in Sec. 4.3: moderate truncation effectively filters the most problematic cases while preserving sufficient coupling diversity for training.
Important Implementation Details.
Table 3: Ablation study on ImageNet-$512^{2}$ for key implementation details of Re-Meanflow.
@@
Base (Best Setting reported in [11]) 7.81
@@
(d) + Avoid high-variance $(t,r)$ region 3.50
@@
To identify the components most critical for stable and high-quality training for Re-Meanflow, we conduct ablations on ImageNet-$512^{2}$ and showcase their impact on FID in Table 3.
(a) Hyperparameter adjustments. Because rectified trajectories are significantly straighter, we reduce the normalization strength in the MeanFlow adaptive loss from $1.0$ to $0.5$ (similar to a Pseudo-Huber loss). We also find that sampling the classifier-free guidance scale from a uniform distribution improves generation quality.
(b) Following Sabour et al. [37], Lu and Song [29], we replace the EDM2 time embedding $\mathrm{emb}(\log\sigma_{t})$ with $\mathrm{emb}(t)$ to obtain more stable Jacobian–vector product computations. We perform a brief alignment stage that trains the new embedding to match the original outputs at the corresponding noise levels. Details of this alignment are provided in Appendix B.1.
(c) Since rectified trajectories are smoother and require less mid-range emphasis, we adopt the U-shaped $t$ distribution from Lee et al. [24] instead of using the standard lognormal schedule.
(d) Re-Meanflow exhibits unusually high variance when $t>0.95$ and $r<0.4$, a phenomenon we analyze in more detail in the Appendix. Inspired by the truncation strategy used in TCM [25], we mitigate this issue by avoding this high-variance time region. This modification not only improves FID, but also accelerates convergence. More detail is presented in Appendix B.3.
(e) Applying our trajectory-distance truncation heuristic (Sec. 4.3) further improves efficiency and generation quality. Additional analysis of distance–error correlations across resolutions, which motivates our truncation threshold, is provided in AppendixC.2.
@@ 6 Conclusion
We presented Re-Meanflow, a simple yet efficient framework that combines the complementary strengths of Rectified Flow [28] and MeanFlow [11]. By leveraging the synergy between trajectory straightening and mean-velocity learning, Re-Meanflow outperforms both components when used independently. Empirically, on ImageNet $64^{2}$, $256^{2}$, and $512^{2}$, Re-Meanflow delivers substantial improvements in both efficiency and generation quality, surpassing current state-of-the-art one-step flow-based models.
Limitation. Since our method does not access real data during training and relies entirely on synthetic samples generated by pretrained diffusion or flow models, its performance naturally depends on the quality of these generated couplings. A promising direction is to incorporate real data into the training process, as explored in Lee et al. [24], Seong et al. [39]. Another complementary line of work investigates how to improve generative models using only their own outputs [2, 3, 4], which suggests that self-generated data can still be leveraged effectively with appropriate regularization. Alternatively, one could improve the synthetic supervision directly by using stronger backbone models to generate the couplings.
Future Work. Looking ahead, the core idea behind Re-Meanflow is architecture-agnostic. We believe that leveraging straighter trajectories is a broadly applicable principle that may benefit other few-step generative paradigms. Re-Meanflow also offers practical advantages by shifting most of the training burden from scarce high-end training GPUs to widely available inference-grade accelerators capable of performing the reflow process. As a proof of concept, we view Re-Meanflow as a promising direction to scaling one-step paradigms to larger domains, such as text-to-image generation.
@@
Appendix A More Discussion
In this section, we provide additional analysis and context for our method. We begin by discussing the broader impact of our approach. We then revisit the theoretical motivation for using a single reflow iteration and explain how our perspective complements existing analyses. Finally, we compare our method with the concurrent CMT method [19].
A.1 Broader Impact
This work introduces a simple yet effective framework for efficient one-step generative modeling by training a MeanFlow model on rectified trajectories. Through extensive evaluations across multiple ImageNet settings, varying in resolution, latent representation, and model architecture, we demonstrate that Re-Meanflow achieves strong performance and high efficiency.
Beyond empirical gains, our method highlights a practical and socially relevant shift in how large generative models may be trained. Traditional few-step or one-step distillation pipelines often require expensive, high-end training hardware (e.g., A100-class GPUs), limiting accessibility to well-resourced institutions. In contrast, Re-Meanflow moves many of the computational burden to the inference reflow stage, followed by a lightweight MeanFlow training phase. Because inference workloads can be executed efficiently on widely available consumer- or inference-grade accelerators, our framework reduces reliance on scarce training GPUs and lowers the barrier to experimentation.
A.2 Additional Details for Sec. 4.1
Lee et al. [24] analyze when multiple Reflow iterations are truly necessary for achieving straight trajectories in rectified flows. Their argument centers on how trajectory intersections affect the learned velocity field.
Consider two 1-rectified couplings $(\mathbf{x}^{\prime},\mathbf{z}^{\prime})$ and $(\mathbf{x}^{\prime\prime},\mathbf{z}^{\prime\prime})$. A trajectory intersection occurs if there exists $t\in[0,1]$ such that
@@ A.1 Broader Impact
To understand how often this phenomenon can happen, Lee et al. [24] show that an intersection implies
@@
Under typical training, nearly all noise samples used to form 1-rectified couplings lie in the high-density region of the Gaussian prior. The $\mathbf{z}^{\prime\prime}$ required above usually lies far outside that region unless $\mathbf{x}^{\prime}-\mathbf{x}^{\prime\prime}||_{2}$ is extremely small or $t$ is very close to 1. Therefore, intersections are statistically rare.
Further assuming that the 1-rectified flow is approximately $L$-Lipschitz,
@@
Relation to Our Work.
Our focus is complementary to the above perspective. While their analysis shows that intersections are rare for most couplings, our experiments reveal that in realistic settings, a small but influential subset of couplings can still exhibit noticeable curvature, particularly when the effective Lipschitz constant $L$ becomes large due to geometric imbalance in the data distribution. Our method is designed to improve robustness in these challenging cases, providing stable one-step generation even when residual curvature persists after a single reflow step.
A.3 Comparison with CMT [19]
CMT [19] is an important concurrent effort that also leverages synthetic trajectories generated by a pretrained sampler to stabilize few-step flow-map training. Conceptually, CMT introduces a dedicated mid-training stage that learns a full trajectory-to-endpoint mapping from solver-generated paths, which then serves as a trajectory-aligned initialization for a subsequent post-training flow-map stage. In contrast, our method adopts a fundamentally different design: rather than supervising on entire solver trajectories, we distill only the end-point couplings of rectified flows and learn the corresponding mean velocity in a single training stage. This distinction yields a practical advantage—our pipeline avoids the compounded complexity of CMT’s two-stage optimization, which is sensitive to hyperparameters at both stages and substantially more expensive to tune.
Appendix B Implementation Details
@@
Model backbone EDM2-S[23] SiT-XL[31] EDM2-S[23]
@@
Learning rate $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
@@
Avoiding high-variance $(t,r)$ $t>0.95\wedge r<0.4$ $t>0.95\wedge r<0.4$ $t>0.95\wedge r<0.4$
Sampling details (rectified couplings from 1-rectified flow)
Pretrained model EDM2-S[23] SiT-XL[31] EDM2-S[23]
@@
Guidance method Autoguidance[22] CFG[16] Autoguidance[22]
@@
B.1 Pretrained Model Conditioning
@@
On ImageNet $64^{2}$ and $512^{2}$, which Re-Meanflow is initialized Re-Meanflow from the pretrained EDM2-S model [23], following AYS [37], we perform a short alignment stage in which we train the new embeddings to reproduce the original time-embedding output for the corresponding noise level. Specifically, for EDM2-S the original embedding depends on $\log\sigma_{t}$, where $\sigma_{t}=\tfrac{t}{1-t}$. We train the new embeddings via:
@@
for 10k iterations with learning rate 1e-3. This process takes only a few minutes. We also convert the original VE diffusion parameterization into the flow-matching setting following Lee et al. [24].
On ImageNet $256^{2}$, we initialize Re-Meanflow from SiT-XL [31], which is already a flow-based model. In this case, only the additional r-embedding needs to be introduced. We simply zero-initialize $\mathrm{emb}_{r}(r)$ and keep the original SiT time embedding for $\mathrm{emb}_{t}(t)$.
B.2 Time Distribution
We observe a similar loss–time profile to that reported in Lee et al. [24]: the training loss as a function of $t$ closely matches that of 2-rectified flow++. Accordingly, for sampling $t$ we adopt the same U-shaped distribution:
@@ B.2 Time Distribution
Following AYF [37], after sampling $t$ we draw the interval length |tr||t-r| from a normal distribution $\mathcal{N}(P_{\text{mean}},P_{\text{std}})$ and apply a sigmoid transformation. We use the same parameters as AYF, $(P_{\text{mean}},P_{\text{std}})=(-0.8,1.0)$, which emphasize medium-length intervals and substantially improve stability. As shown in Table 3, this setting yields a noticeable improvement in FID and accelerates convergence.
We also experimented with sampling $t$ uniformly. Despite its simplicity, uniform sampling performs competitively, and in many cases better than commonly used log-normal time distributions in diffusion and flow-matching models [21, 23, 11], which prioritize mid-range timesteps to avoid high variance near $t\approx 0$ and $t\approx 1$. We attribute the strong performance of uniform sampling to the significantly lower variance of rectified trajectories: after one reflow step, early-time and high-noise regions become much more stable, allowing us to allocate more samples to these challenging regimes without the usual degradation observed when training on independent couplings.
B.3 Avoiding High-Variance Time Regions
@@
Inspired by the truncation strategy in TCM [25], we adopt a simple yet effective rule to avoid this problematic regime: whenever a sampled pair satisfies $t>0.95$ and $r<0.4$, we set $r=0$. Empirically, this improves both training stability and FID. Our hypothesis is that predicting a clean image from pure noise ($r=0$, $t\approx 1$) is substantially easier than predicting a lightly corrupted target: the latter requires the model to determine which noise components should be preserved, introducing ambiguity and variance at high $t$. By redirecting training toward these easier high-$t$ targets, the model can allocate more capacity to learning accurate one-step predictions. This modification not only improves FID relative to configuration (c), but also accelerates convergence: in the high-$t$ regime, more updates involve $r=0$, allowing the model to refine its one-step outputs more quickly and reliably.
B.4 Training with Guidance
Classifier-free guidance (CFG) [16] is widely used to boost the performance of diffusion and flow-based generative models. To incorporate CFG into the MeanFlow stage, we train Re-Meanflow on the CFG-enhanced velocity field:

### v2

@@ 4.3 Loss Landscape
Re-MeanFlow (top $15\%$) 3.19 259.61 0.76 0.53
ImageNet $256^{2}$
MeanFlow [geng2025mean] 3.43 247.5 0.73 0.54
Re-MeanFlow (ours; no trunc.) 3.48 243.4 0.72 0.54
[…]
4.4 Distance-based Truncation
As discussed in Sec. 3.4.2, we empirically observe that large endpoint distances $\|\mathbf{z}-\mathbf{x}\|_{2}$ are correlated with higher trajectory curvature. We therefore apply a simple top-$10\%$ truncation rule, discarding the couplings with the largest $\ell_{2}$ endpoint distance. In practice, this filter removes many residual high-curvature pairs, improving training stability and sample quality for Re-MeanFlow.
To study the effect of the truncation ratio, we sweep different truncation rates and report results on ImageNet $512^{2}$ and ImageNet $256^{2}$ (Tab. 2). We find that moderate truncation consistently improves stability and yields better IS/FID, while maintaining diversity as reflected by comparable Recall. In particular, FID improves monotonically within a moderate truncation regime.
Interestingly, even without truncation, Re-MeanFlow already matches the performance of MeanFlow trained directly on real data. This is notable because Re-MeanFlow does not require access to the original dataset: it learns solely from generated coupling pairs. Since self-generated distillation typically incurs discretization error that can degrade performance, we hypothesize that the rectified trajectories substantially simplify the optimization problem, enabling more effective parameter learning and thus competitive (or improved) results.
@@ 4.5 Ablation Study
Table 3: Ablation study on ImageNet-$512^{2}$ for key implementation details.
@@
Base (Best Setting reported in [geng2025mean]) 7.81
@@
(d) + Avoid high-variance $(r,t)$ region 3.50
@@
To identify the implementation choices most critical for stable and high-quality training, we conduct an ablation study on ImageNet-$512^{2}$ in Tab. 3. We summarize the main findings below.
(a) Hyperparameter adjustments. Since rectified trajectories are substantially straighter, we reduce the normalization strength in MeanFlow’s adaptive loss from $1.0$ to $0.5$ (Pseudo-Huber style).
(b) Time embedding change. Following [sabour2025align, lu2024simplifying], we replace the EDM2 time embedding $\mathrm{emb}(\log\sigma_{t})$ with $\mathrm{emb}(t)$ to improve the stability of Jacobian-vector product computations.
(c) U-shaped $t$ distribution. Because rectified trajectories reduce the need for mid-range emphasis, we adopt the U-shaped $t$ distribution from [lee2024improving] in place of the standard lognormal schedule.
(d) Avoid high-variance $(r,t)$ region. We observe unusually high variance when $t>0.95$ and $r<0.4$ (analyzed in the Appendix  0.B.3). Inspired by the truncation strategy in TCM [lee2024truncated], we exclude this time region, which improves FID and accelerates convergence.
(e) Distance-based truncation. Finally, applying our distance-based truncation heuristic (Sec. 3.4.2) further improves both efficiency and generation quality.
@@ 5 Conclusion
We introduced Rectified-MeanFlow, a lightweight, data-free self-distillation framework for one-step generation that addresses a key MeanFlow bottleneck: learning mean-velocity on highly curved trajectories. Our central geometric takeaway is that mean-velocity estimation is substantially easier on straighter paths. Accordingly, Re-MeanFlow models mean-velocity on rectified couplings generated from a pretrained flow model, yielding a simpler velocity field and a markedly better-conditioned loss landscape; this translates into faster convergence and stronger one-step generation. We further improve robustness with a distance-based truncation rule that removes residual high-curvature couplings. Empirically, on ImageNet at $64^{2}$, $256^{2}$, and $512^{2}$, Re-MeanFlow consistently outperforms prior one-step flow distillation methods and strong Rectified Flow baselines, achieving higher sample quality with substantially less training.
@@
Appendix 0.A More Discussion
In this section, we provide additional analysis and context for our method. We begin by discussing its limitations and then examine the broader impact of our approach. Next, we revisit the theoretical motivation for using a single rectification iteration and explain how our perspective complements existing analyses. Finally, we compare our method with the concurrent CMT method [hu2025cmt].
0.A.1 Limitation
Since our method does not access real data during training and relies entirely on synthetic samples generated by pretrained diffusion or flow models, its performance naturally depends on the quality of these generated couplings. A promising direction is to incorporate real data into the training process, as explored in [lee2024improving, seong2025balanced]. Another complementary line of work investigates how to improve generative models using only their own outputs [alemohammad2023self, alemohammad2024self, alemohammad2025neon], which suggests that self-generated data can still be leveraged effectively with appropriate regularization. Alternatively, one could improve the synthetic supervision directly by using stronger backbone models to generate the couplings.
0.A.2 Broader Impact
Beyond empirical gains, our method highlights a practical and socially relevant shift in how large generative models can be trained. Traditional few-step or one-step distillation pipelines often concentrate compute in expensive training workloads on high-end accelerators (e.g., A100-class GPUs), which limits accessibility to well-resourced institutions. In contrast, Re-MeanFlow shifts a substantial fraction of computation to an inference-only stage (sampling rectified couplings), followed by a lightweight MeanFlow training phase. Because inference workloads can be executed efficiently on widely available consumer- or inference-grade accelerators, our framework reduces reliance on scarce training GPUs and lowers the barrier to experimentation.
Practically, the sampling process for the rectified couplings is amenable to pipeline parallelism: coupling generation can run asynchronously alongside training as a data-preparation stream. For implementation convenience, we pre-generate couplings before training in this work; nevertheless, even under this conservative setup, Re-MeanFlow achieves the lowest total compute among prior state-of-the-art distillation methods, underscoring the end-to-end efficiency of our framework.
0.A.3 Comparison with One-Step Sampling via Rectified Flow
Rectified Flow [liu2022flow] supports few-step sampling via an iterative reflow procedure that progressively straightens trajectories by repeatedly training a new flow model on rectified couplings. Liu et al. [liu2022flow] show that, in the limit, iterating this procedure can yield perfectly straight couplings. In practice, however, achieving near-linear paths for reliable one-step sampling typically requires multiple reflow rounds, which is computationally expensive and can degrade performance due to accumulated approximation errors.
Lee et al. [lee2024improving] analyze when multiple Reflow iterations are truly necessary for achieving straight trajectories in rectified flows. Their argument centers on how trajectory intersections affect the learned velocity field. Consider two 1-rectified couplings $(\mathbf{x}^{\prime},\mathbf{z}^{\prime})$ and $(\mathbf{x}^{\prime\prime},\mathbf{z}^{\prime\prime})$. A trajectory intersection occurs if there exists $t\in[0,1]$ such that
@@ 0.A.3 Comparison with One-Step Sampling via Rectified Flow
To understand how often this phenomenon can happen, Lee et al. [lee2024improving] show that an intersection implies
@@
Under typical training, nearly all noise samples used to form 1-rectified couplings lie in the high-density region of the Gaussian prior. The $\mathbf{z}^{\prime\prime}$ required above usually lies far outside that region unless $\mathbf{x}^{\prime}-\mathbf{x}^{\prime\prime}||_{2}$ is extremely small or $t$ is very close to 1. Therefore, intersections are statistically rare. Further assuming that the 1-rectified flow is approximately $L$-Lipschitz,
@@
0.A.3.1 Relation to Our Work.
Our focus is complementary. While the above analysis suggests that trajectory intersections are rare for most couplings, our experiments indicate that realistic settings can still contain a small but influential subset of pairs with non-negligible curvature, particularly when the effective Lipschitz constant $L$ is large due to geometric imbalance in the data distribution. As shown in Fig. 2b and Fig. 11, trajectories induced by a once-rectified coupling (i.e., one reflow step) are substantially straighter than those from the independent coupling (Fig. 2a), yet residual curvature remains and can still hinder reliable one-step sampling based on the instantaneous velocity. Re-MeanFlow is designed to be robust in precisely these challenging cases, enabling stable one-step generation even when curvature persists after a single rectification step.
0.A.4 Comparison with CMT
CMT [hu2025cmt] is an important concurrent effort that also leverages synthetic trajectories generated by a pretrained sampler to stabilize few-step flow-map training. Conceptually, CMT introduces a dedicated mid-training stage that learns a full trajectory-to-endpoint mapping from solver-generated paths, which then serves as a trajectory-aligned initialization for a subsequent post-training flow-map stage. In contrast, our method adopts a fundamentally different design: rather than supervising on entire solver trajectories, we distill only the end-point couplings of rectified flows and learn the corresponding mean velocity in a single training stage. This distinction yields a practical advantage: our pipeline avoids the compounded complexity of CMT’s two-stage optimization, which is sensitive to hyperparameters at both stages and substantially more expensive to tune.
Appendix 0.B Implementation Details
@@
Model Backbone EDM2-S[karras2024analyzing] SIT-XL[ma2024sit] EDM2-S[karras2024analyzing]
@@
Learning Rate 1e-4 1e-4 1e-4
@@
Avoiding High-variance $(t,r)$ Region t>0.95&r<0.4t>0.95\ \&\ r<0.4 t>0.95&r<0.4t>0.95\ \&\ r<0.4 t>0.95&r<0.4t>0.95\ \&\ r<0.4
Sampling Details for Rectified Couplings
Pretrained Model EDM2-S[karras2024analyzing] SIT-XL[ma2024sit] EDM2-S[karras2024analyzing]
@@
Guidance Method Autoguidance [karras2024guiding] CFG [ho2022classifier] Autoguidance [karras2024guiding]
@@
0.B.1 Pretrained Model Conditioning
@@
0.B.1.1 On ImageNet $64^{2}$ and $512^{2}$
, which Re-MeanFlow is initialized from the pretrained EDM2-S model [karras2024analyzing], following AYF [sabour2025align], we perform a short alignment stage in which we train the new embeddings to reproduce the original time-embedding output for the corresponding noise level. Specifically, for EDM2-S the original embedding depends on $\log\sigma_{t}$, where $\sigma_{t}=\tfrac{t}{1-t}$. We train the new embeddings via:
@@ 0.B.1.1 On ImageNet $64^{2}$ and $512^{2}$
for 10k iterations with learning rate 1e-3. This process takes only a few minutes. We also convert the original VE diffusion parameterization into the flow-matching setting following [lee2024improving].
0.B.1.2 On ImageNet $256^{2}$
, we initialize Re-MeanFlow from SiT-XL [ma2024sit], which is already a flow-based model. In this case, only the additional r-embedding needs to be introduced. We simply zero-initialize $\mathrm{emb}_{r}(r)$ and keep the original SiT time embedding for $\mathrm{emb}_{t}(t)$.
0.B.2 Time Distribution
We observe a similar loss–time profile to that reported in [lee2024improving]: the training loss as a function of $t$ closely matches that of 2-rectified flow++. Accordingly, for sampling $t$ we adopt the same U-shaped distribution:
@@ 0.B.2 Time Distribution
Following AYF [sabour2025align], after sampling $t$ we draw the interval length |tr||t-r| from a normal distribution $\mathcal{N}(P_{\text{mean}},P_{\text{std}})$ and apply a sigmoid transformation. We use the same parameters as AYF, $(P_{\text{mean}},P_{\text{std}})=(-0.8,1.0)$, which emphasize medium-length intervals and substantially improve stability. As shown in Table 3, this setting yields a noticeable improvement in FID and accelerates convergence.
We also experimented with sampling $t$ uniformly. Despite its simplicity, uniform sampling performs competitively—and in many cases better—than commonly used log-normal time distributions in diffusion and flow-matching models [karras2022elucidating, karras2024analyzing, geng2025mean], which prioritize mid-range timesteps to avoid high variance near $t\approx 0$ and $t\approx 1$. We attribute the strong performance of uniform sampling to the significantly lower variance of rectified trajectories: after one reflow step, early-time and high-noise regions become much more stable, allowing us to allocate more samples to these challenging regimes without the usual degradation observed when training on independent couplings.
@@ 0.B.3 Avoiding High-Variance Time Regions
Inspired by the truncation strategy in TCM [lee2024truncated], we adopt a simple yet effective rule to avoid this problematic regime: whenever a sampled pair satisfies $t>0.95$ and $r<0.4$, we set $r=0$. Empirically, this improves both training stability and FID. Our hypothesis is that predicting a clean image from pure noise ($r=0$, $t\approx 1$) is substantially easier than predicting a lightly corrupted target: the latter requires the model to determine which noise components should be preserved, introducing ambiguity and variance at high $t$. By redirecting training toward these easier high-$t$ targets, the model can allocate more capacity to learning accurate one-step predictions. This modification not only improves FID relative to configuration (c), but also accelerates convergence: in the high-$t$ regime, more updates involve $r=0$, allowing the model to refine its one-step outputs more quickly and reliably.
0.B.4 Training with Guidance
Classifier-free guidance (CFG) [ho2022classifier] is widely used to boost the performance of diffusion and flow-based generative models. To incorporate CFG into the MeanFlow stage, we train Re-MeanFlow on the CFG-enhanced velocity field: